# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v39

@@ 6.3 Example - Binary Signal Case
Suppose there are two possible information stories, $S=\{s_{1},s_{2}\}$. In the canonical game there is one contract per story, paying $1$ if that story is true and 0 otherwise. The natural informed trade is a market-neutral spread: under each signal the insider buys the “active” contract and shorts the alternative by the same amount. This long-short trade corresponds to the strategy matrix
@@
In this two-story case, the cross-section of information-contract prices $(p_{1},p_{2})$ is fully summarized by a single posterior log-odds statistic $Z$:
@@
Thus observing order flow is equivalent to observing the “evidence score” $Z$: large $Z$ is strong evidence for $s_{1}$, while large negative $Z$ favors $s_{2}$. [^18] See Lemma A.2 in the Appendix.
@@
This tradeoff is characterized by the marginal indifference condition for $\alpha$ (our scalar equilibrium equation) [^19] See Equation (A.23) in the Appendix.
@@
where the expectations are taken under (27) (and hence are functions of $\alpha$). Here $1-\mathbb{E}[p_{1}]$ is the average “wedge” the insider earns from the active story, while $p_{1}p_{2}=p_{1}(1-p_{1})$ is the posterior variance: it is largest when $Z\approx 0$ (the market is unsure) and small when $|Z|$ is large (the market is confident), making precise how marginal scaling of the spread further sharpens inference and erodes informational rents.
@@ 6.4 Comparative Statics of $\alpha^{*}$
$S=\{s_{1},\ldots,s_{K}\}$ Set of possible information stories (signals) about the payoff distribution (e.g., volatility, tail risk, skewness).
@@
$\Lambda=(\Lambda_{i,j})$ Equilibrium price-impact matrix: $\Lambda_{i,j}$ is the response of $P^{*}(x_{i}\mid\omega)$ to order flow in claim $x_{j}$ (own-impact on diagonals, cross-impact off-diagonals).
@@ 7 Price Discovery
For ease of reference, Table 1 recollects the key equilibrium quantities, organized by our model’s information channel:
@@
The private signal $s_{k}$ induces equilibrium informed demand $W^{*}(\,\cdot\,\mid s_{k})$, which is part of aggregate order flow $\omega$. Conditioning on $\omega$, the market maker sets Arrow-Debreu state prices $P^{*}(\,\cdot\,\mid\omega)$; the within- and cross-price responses are collected in the price impact matrix $\Lambda$. Applying the standard formula (7) expresses these Arrow–Debreu quantities in terms of options.
@@ 7.2 Price Impact
We next characterize the equilibrium price impact matrix $\Lambda$. Its $(i,j)$-element $\Lambda_{i,j}$ is the equilibrium marginal effect of trades in $x_{j}$ on the price of $x_{i}$ (diagonals are within-market impact; off-diagonals are cross impact). The reader can find the notations used in this subsection recalled in Table 1.
@@
  • Informational linkage: The remaining quantity,
@@
    • $\circ$
      Toy model illustration: we already saw this mechanism in Section 2—when two claims load on the same underlying risk (e.g., the put-call legs of a straddle under a volatility signal), their prices move together. The next example further illustrates this price impact sign pattern.
@@ 7.3 Information Efficiency of Prices
Our contribution is complementary and structural: conditional on aggregate order flow $\omega$, equilibrium prices embed a posterior over signals, $\pi_{1}^{*}(\,\cdot\,\mid\omega)$ (see Table 1), which summarizes the informativeness of the entire cross-section of prices. We therefore begin by characterizing this posterior in closed form.
@@ 8 Empirical Implications
Although our framework applies to general contingent-claim markets, we discuss empirical implications through the lens of equity options. Options are the natural setting in which information beyond the expected payoff can be traded, and the market for which rich price and order flow data have supported extensive empirical work.
@@
In characterizing the equilibrium informational channel, we map distinct informational dimensions map to distinct option portfolios, and they load into the option surface through systematic cross-impact patterns across strikes and strategies. This mapping is not restricted to higher-moment information: it applies to any feature spanned by option payoffs, yielding feature-specific predictions for sign and concentration patterns in $\widehat{\Lambda}$ across strikes and strategies.
@@ 9 Conclusion
Building on Arrow and Debreu (1954) and Kyle (1985), we develop a general, tractable model of price discovery across contingent-claim markets. By rationalizing observed market practices in an equilibrium setting, the model helps bridge a longstanding gap between empirical regularities and theory. Its tractability and generality invite natural follow-on questions. How does risk aversion reshape informed portfolios and the pricing kernel—and how does asymmetric information distort risk sharing across contingent claims? In a dynamic setting, how do information spillovers propagate across strikes and maturities? Beyond equity options, the same approach can be applied to futures and forwards, credit derivatives, and interest-rate derivatives, where order-flow-driven prices map into term-structure, default-intensity, and correlation information. These and related extensions provide a natural agenda for studying information transmission across diverse derivative markets.

### v40

@@ 6.3 Example - Binary Signal Case
Suppose there are two possible information stories, $S=\{s_{1},s_{2}\}$. In the canonical game there is one information contract per story, paying $1$ if that story is true and 0 otherwise. The natural informed trade is a market-neutral spread: under each signal the insider buys the “active” contract and shorts the alternative by the same amount. This long-short trade corresponds to the strategy matrix
@@
In this two-story case, the cross-section of information-contract prices $(p_{1},p_{2})$ is fully summarized by a single posterior log-odds statistic ZZ: [^18] This is a special case of Lemma A.2.
@@
Thus observing order flow is equivalent to observing the “evidence score” $Z$: large $Z$ is strong evidence for $s_{1}$, while large negative $Z$ favors $s_{2}$. In the toy setting of Section 2, the evidence score reduces to the normalized cross-market imbalance $\Delta(\omega)$ in (2), with $Z=2\alpha\,\Delta(\omega)$. Accordingly, in the volatility interpretation of Section 2, large $Z$ is strong evidence for high volatility.
@@
Our scalar equilibrium equation is exactly the marginal indifference condition for the trading scale $\alpha$:
@@
where the expectations are taken under (27) (and hence are functions of $\alpha$). [^19] This is an instance of the general equilibrium equation (A.23). Here $1-\mathbb{E}[p_{1}]$ is the average “wedge” the insider earns from the active story, while $p_{1}p_{2}=p_{1}(1-p_{1})$ is the posterior variance: it is largest when $Z\approx 0$ (the market is unsure) and small when $|Z|$ is large (the market is confident), making precise how marginal scaling of the spread further sharpens inference and erodes marginal information rents.
@@ 6.4 Comparative Statics of $\alpha^{*}$
$S=\{s_{1},\ldots,s_{K}\}$ Set of possible information stories (signals) about the payoff distribution (e.g., volatility, skewness, tail risk).
@@
$\Lambda=(\Lambda_{i,j})$ Equilibrium price impact matrix: $\Lambda_{i,j}$ is the response of $P^{*}(x_{i}\mid\omega)$ to order flow in claim $x_{j}$ (own-impact on diagonals, cross-impact off-diagonals).
@@ 7 Price Discovery
For ease of reference, Table LABEL:tab:key_equilibrium_quantities recollects the key equilibrium quantities, organized by the model’s information channel:
@@
The private signal $s_{k}$ induces equilibrium informed demand $W^{*}(\,\cdot\,\mid s_{k})$, which is part of aggregate order flow $\omega$. Conditioning on $\omega$, the market maker sets Arrow-Debreu state prices $P^{*}(\,\cdot\,\mid\omega)$; the within- and cross-price responses are collected in the price impact matrix $\Lambda$. The standard formula (7) translates these Arrow–Debreu quantities to options.
@@ 7.2 Price Impact
We next characterize the equilibrium price impact matrix $\Lambda$. Its $(i,j)$-element $\Lambda_{i,j}$ is the equilibrium marginal effect of trades in $x_{j}$ on the price of $x_{i}$ (diagonals are within-market impact; off-diagonals are cross impact). The reader can find the notations used in this subsection recalled in Table LABEL:tab:key_equilibrium_quantities.
@@
  • Information linkage: The remaining quantity,
@@
We already saw the information-linkage intuition expressed directly through payoff covariances in the toy model of Section 2—when two claims load on the same underlying risk (e.g., the put–call legs of a straddle under a volatility signal), their prices move in tandem. Having established that this intuition carries over to the general setting, the next example provides a further concrete illustration.
@@ 7.3 Information Efficiency of Prices
Our contribution is complementary and structural: conditional on aggregate order flow $\omega$, equilibrium prices embed a posterior over signals, $\pi_{1}^{*}(\,\cdot\,\mid\omega)$ (see Table LABEL:tab:key_equilibrium_quantities), which summarizes the informativeness of the entire cross-section of prices. We therefore begin by characterizing this posterior in closed form.
@@ 8 Empirical Implications
While our framework applies to general contingent-claim markets, we examine its immediate empirical implications through the lens of equity options. Options are the natural setting in which information beyond the expected payoff can be traded, and the market for which rich price and order flow data have supported extensive empirical work.
@@
In characterizing the equilibrium informational channel, we map distinct informational dimensions to distinct option portfolios, and they load into the option surface through systematic cross-impact patterns across strikes and strategies. This mapping is not restricted to higher-moment information: it applies to any feature spanned by option payoffs, yielding feature-specific predictions for sign and concentration patterns in $\widehat{\Lambda}$ across strikes and strategies.
@@ 9 Conclusion
Building on Arrow and Debreu (1954) and Kyle (1985), we develop a general model of price discovery across contingent-claim markets. In the same spirit as Kyle (1985)—which rationalizes intuitive informed-trading behavior in the single-asset case—our Arrow–Debreu generalization elevates that equilibrium rationale to settings in which many contingent claims trade jointly. The result is a constructive and tractable equilibrium account of observed market practices and empirical regularities across derivative markets, along with new, testable predictions.
The tractability and generality of our model invite natural follow-on questions. How does risk aversion reshape informed portfolios and the pricing kernel, and how does asymmetric information distort risk sharing across contingent claims? In a dynamic setting, how do information spillovers propagate across strikes and maturities? Beyond equity options, the same approach can be applied to futures and forwards, credit derivatives, and interest-rate derivatives, where order-flow-driven prices map into term-structure, default-intensity, and correlation information. We hope to pursue these and related questions in future work.